OpenStudy (anonymous):

An electric motor has a power rating of 1.50 kW. If it operates at 70% efficiency, what work can it do in an hour?

OpenStudy (mrnood):

a) Do you know the relationship between power and work? b) If the rated power is 1.5 kW - what is the effective output power at 70% efficiency. Post your answers here - I'll help more if I'm around...

OpenStudy (anonymous):

do u go P=w/t 1125W/3600s=0.3125/w/s? @MrNood

OpenStudy (mrnood):

You are on the right track. However - firstly - you have used 75% not 70% (so the effective power is 1050W not 1125W) Second P=w/t - so what is the correct equation for w=? Finally - the units are not correct (but this will be corrected when you get the equation right...) You're nearly there -

OpenStudy (anonymous):

so W=P*T 1050W*3600s=3.8x10^6W/s? im kinda confused about the units @mrnood

OpenStudy (mrnood):

You have been asked to find WORK - its units are Joules. You have multiplied Power*t - BUT you have written W/s It should be Wxs I W = INm/s SO 1 W x !s = 1Nm = I J You have the numbers correct - but the units are equally important

OpenStudy (mrnood):

Also - why have you rounded the answer to 3.8? It is 3.78. The power was given to 3 sf - you should keep the same precision